# LaT-PFN: A Joint Embedding Predictive Architecture for In-context Time-series Forecasting

### v1

@@ Appendix D Latent Space Visualised
Appendix E Additional results
E.1 UCR [20] Accuracy by Dataset
@@
ACSF1_TEST 37.00% 41.00%
Adiac_TEST 2.05% 2.05%
AllGestureWiimoteX_TEST 29.14% 25.14%
AllGestureWiimoteY_TEST 25.00% 17.86%
AllGestureWiimoteZ_TEST 22.86% 16.43%
ArrowHead_TEST 46.29% 41.71%
BME_TEST 54.00% 46.67%
Beef_TEST 50.00% 40.00%
BeetleFly_TEST 60.00% 85.00%
BirdChicken_TEST 60.00% 75.00%
CBF_TEST 53.33% 47.56%
Car_TEST 21.67% 21.67%
Chinatown_TEST 60.06% 76.38%
ChlorineConcentration_TEST 53.26% 53.26%
CinCECGTorso_TEST 29.06% 25.58%
Coffee_TEST 75.00% 82.14%
Computers_TEST 56.40% 66.40%
CricketX_TEST 29.23% 13.85%
CricketY_TEST 24.87% 17.44%
CricketZ_TEST 29.49% 14.36%
Crop_TEST 38.66% 32.28%
DiatomSizeReduction_TEST 30.07% 30.07%
DistalPhalanxOutlineAgeGroup_TEST 67.63% 46.76%
DistalPhalanxOutlineCorrect_TEST 58.33% 58.33%
DistalPhalanxTW_TEST 58.27% 30.22%
DodgerLoopDay_TEST 23.75% 37.50%
DodgerLoopGame_TEST 56.52% 50.00%
DodgerLoopWeekend_TEST 87.68% 89.13%
ECG200_TEST 64.00% 64.00%
ECG5000_TEST 82.69% 87.51%
ECGFiveDays_TEST 49.71% 49.71%
EOGHorizontalSignal_TEST 17.13% 29.83%
EOGVerticalSignal_TEST 21.82% 20.99%
Earthquakes_TEST 74.82% 74.82%
ElectricDevices_TEST 51.87% 54.30%
EthanolLevel_TEST 25.40% 26.40%
FaceAll_TEST 26.63% 42.43%
FaceFour_TEST 44.32% 18.18%
FacesUCR_TEST 27.46% 18.29%
FiftyWords_TEST 22.64% 12.53%
Fish_TEST 14.29% 22.29%
FordA_TEST 76.52% 61.52%
FordB_TEST 65.43% 53.95%
FreezerRegularTrain_TEST 76.70% 75.75%
FreezerSmallTrain_TEST 76.00% 75.86%
Fungi_TEST 60.22% 46.24%
@@
GestureMidAirD1_TEST 20.77% 20.00%
GestureMidAirD2_TEST 26.92% 20.77%
GestureMidAirD3_TEST 16.15% 10.00%
GesturePebbleZ1_TEST 40.12% 50.58%
GesturePebbleZ2_TEST 42.41% 31.01%
GunPointAgeSpan_TEST 58.23% 57.91%
GunPointMaleVersusFemale_TEST 62.66% 84.81%
GunPointOldVersusYoung_TEST 52.38% 100.00%
GunPoint_TEST 49.33% 64.67%
Ham_TEST 59.05% 51.43%
HandOutlines_TEST 64.05% 64.05%
Haptics_TEST 22.40% 20.78%
Herring_TEST 59.38% 59.38%
HouseTwenty_TEST 69.75% 71.43%
InlineSkate_TEST 18.18% 18.73%
InsectEPGRegularTrain_TEST 100.00% 100.00%
InsectEPGSmallTrain_TEST 100.00% 100.00%
InsectWingbeatSound_TEST 30.15% 21.26%
ItalyPowerDemand_TEST 61.22% 49.85%
LargeKitchenAppliances_TEST 56.27% 57.60%
Lightning2_TEST 54.10% 54.10%
Lightning7_TEST 30.14% 41.10%
Mallat_TEST 25.54% 12.32%
Meat_TEST 68.33% 53.33%
MedicalImages_TEST 51.45% 51.45%
MelbournePedestrian_TEST 24.52% 62.03%
MiddlePhalanxOutlineAgeGroup_TEST 45.45% 18.83%
MiddlePhalanxOutlineCorrect_TEST 57.04% 57.04%
MiddlePhalanxTW_TEST 46.10% 27.27%
MixedShapesRegularTrain_TEST 59.13% 52.21%
MixedShapesSmallTrain_TEST 48.78% 48.49%
MoteStrain_TEST 87.22% 77.32%
NonInvasiveFetalECGThorax1_TEST 21.63% 1.83%
NonInvasiveFetalECGThorax2_TEST 26.97% 1.83%
OSULeaf_TEST 35.95% 18.18%
OliveOil_TEST 40.00% 40.00%
Overall 51.20% 45.74%
PLAID_TEST 31.10% 24.39%
PhalangesOutlinesCorrect_TEST 61.31% 61.31%
Phoneme_TEST 15.98% 11.29%
PickupGestureWiimoteZ_TEST 38.00% 52.00%
PigAirwayPressure_TEST 5.77% 19.71%
PigArtPressure_TEST 4.81% 25.96%
PigCVP_TEST 5.29% 46.63%
Plane_TEST 22.86% 9.52%
PowerCons_TEST 87.22% 85.56%
ProximalPhalanxOutlineAgeGroup_TEST 82.44% 48.78%
ProximalPhalanxOutlineCorrect_TEST 68.38% 68.38%
ProximalPhalanxTW_TEST 63.90% 35.12%
RefrigerationDevices_TEST 39.20% 54.40%
Rock_TEST 46.00% 34.00%
@@
ScreenType_TEST 45.33% 40.80%
SemgHandGenderCh2_TEST 74.33% 58.83%
SemgHandMovementCh2_TEST 30.89% 35.11%
SemgHandSubjectCh2_TEST 37.11% 35.56%
ShakeGestureWiimoteZ_TEST 54.00% 64.00%
ShapeletSim_TEST 50.56% 87.78%
ShapesAll_TEST 28.00% 42.67%
SmallKitchenAppliances_TEST 40.53% 55.47%
SmoothSubspace_TEST 57.33% 78.00%
SonyAIBORobotSurface1_TEST 59.73% 42.93%
SonyAIBORobotSurface2_TEST 78.07% 61.70%
StarLightCurves_TEST 82.93% 84.74%
Strawberry_TEST 64.32% 64.32%
SwedishLeaf_TEST 31.68% 8.48%
Symbols_TEST 55.78% 17.39%
SyntheticControl_TEST 50.33% 77.33%
ToeSegmentation1_TEST 70.61% 65.79%
ToeSegmentation2_TEST 80.77% 80.77%
Trace_TEST 65.00% 19.00%
TwoLeadECG_TEST 49.96% 49.96%
TwoPatterns_TEST 31.62% 25.87%
UMD_TEST 49.31% 50.00%
UWaveGestureLibraryAll_TEST 51.84% 12.37%
UWaveGestureLibraryX_TEST 57.31% 22.28%
UWaveGestureLibraryY_TEST 44.89% 20.38%
UWaveGestureLibraryZ_TEST 52.07% 31.38%
Wafer_TEST 89.21% 89.21%
Wine_TEST 50.00% 50.00%
WordSynonyms_TEST 29.31% 21.94%
WormsTwoClass_TEST 62.34% 57.14%
Worms_TEST 46.75% 42.86%
Yoga_TEST 52.97% 53.57%

### v2

@@ Appendix E Latent Space Visualised
Appendix F Additional results
F.1 UCR [19] Accuracy by Dataset
@@
Acsf1 Test 36.00 % $\pm$ 5.0 % 43.00 % $\pm$ 2.8 %
Adiac Test 03.07 % $\pm$ 2.0 % 02.05 % $\pm$ 0.0 %
Allgesturewiimotex Test 24.50 % $\pm$ 5.8 % 21.21 % $\pm$ 5.6 %
Allgesturewiimotey Test 24.39 % $\pm$ 5.0 % 16.36 % $\pm$ 2.1 %
Allgesturewiimotez Test 19.25 % $\pm$ 4.9 % 17.50 % $\pm$ 1.5 %
Arrowhead Test 43.00 % $\pm$ 10.2 % 44.86 % $\pm$ 4.4 %
Bme Test 51.83 % $\pm$ 2.0 % 49.00 % $\pm$ 3.3 %
Beef Test 41.67 % $\pm$ 8.8 % 38.33 % $\pm$ 2.4 %
Beetlefly Test 56.25 % $\pm$ 4.8 % 75.00 % $\pm$ 14.1 %
Birdchicken Test 67.50 % $\pm$ 15.5 % 70.00 % $\pm$ 7.1 %
Cbf Test 47.92 % $\pm$ 10.8 % 40.33 % $\pm$ 10.2 %
Car Test 21.67 % $\pm$ 0.0 % 21.67 % $\pm$ 0.0 %
Chinatown Test 79.23 % $\pm$ 15.2 % 78.57 % $\pm$ 3.1 %
Chlorineconcentration Test 53.65 % $\pm$ 0.6 % 53.26 % $\pm$ 0.0 %
Cincecgtorso Test 25.85 % $\pm$ 2.1 % 25.18 % $\pm$ 0.6 %
Coffee Test 72.32 % $\pm$ 5.4 % 82.14 % $\pm$ 0.0 %
Computers Test 53.90 % $\pm$ 2.5 % 66.80 % $\pm$ 0.6 %
Cricketx Test 21.15 % $\pm$ 8.7 % 13.85 % $\pm$ 0.0 %
Crickety Test 21.47 % $\pm$ 10.9 % 16.54 % $\pm$ 1.3 %
Cricketz Test 20.45 % $\pm$ 10.4 % 13.85 % $\pm$ 0.7 %
Crop Test 31.78 % $\pm$ 6.2 % 32.07 % $\pm$ 0.3 %
Diatomsizereduction Test 31.70 % $\pm$ 3.3 % 30.07 % $\pm$ 0.0 %
Distalphalanxoutlineagegroup Test 63.49 % $\pm$ 8.3 % 46.76 % $\pm$ 0.0 %
Distalphalanxoutlinecorrect Test 58.33 % $\pm$ 0.0 % 58.33 % $\pm$ 0.0 %
Distalphalanxtw Test 59.35 % $\pm$ 3.7 % 30.22 % $\pm$ 0.0 %
Dodgerloopday Test 26.88 % $\pm$ 5.2 % 38.12 % $\pm$ 0.9 %
Dodgerloopgame Test 60.87 % $\pm$ 11.9 % 48.55 % $\pm$ 2.0 %
Dodgerloopweekend Test 72.10 % $\pm$ 26.6 % 90.22 % $\pm$ 1.5 %
Ecg200 Test 64.00 % $\pm$ 0.0 % 64.00 % $\pm$ 0.0 %
Ecg5000 Test 86.03 % $\pm$ 4.0 % 86.33 % $\pm$ 1.7 %
Ecgfivedays Test 52.41 % $\pm$ 5.4 % 49.71 % $\pm$ 0.0 %
Eoghorizontalsignal Test 16.99 % $\pm$ 2.8 % 29.42 % $\pm$ 0.6 %
Eogverticalsignal Test 16.44 % $\pm$ 5.2 % 21.13 % $\pm$ 0.2 %
Earthquakes Test 74.82 % $\pm$ 0.0 % 74.82 % $\pm$ 0.0 %
Electricdevices Test 52.58 % $\pm$ 3.0 % 52.41 % $\pm$ 2.7 %
Ethanollevel Test 25.95 % $\pm$ 0.6 % 26.30 % $\pm$ 0.1 %
Faceall Test 25.72 % $\pm$ 4.5 % 41.57 % $\pm$ 1.2 %
Facefour Test 38.92 % $\pm$ 10.1 % 27.27 % $\pm$ 12.9 %
Facesucr Test 23.33 % $\pm$ 3.8 % 16.32 % $\pm$ 2.8 %
Fiftywords Test 22.64 % $\pm$ 5.6 % 12.53 % $\pm$ 0.0 %
Fish Test 21.00 % $\pm$ 6.1 % 22.57 % $\pm$ 0.4 %
Forda Test 67.42 % $\pm$ 10.8 % 62.88 % $\pm$ 1.9 %
Fordb Test 57.69 % $\pm$ 6.4 % 56.11 % $\pm$ 3.1 %
Freezerregulartrain Test 76.06 % $\pm$ 0.9 % 75.79 % $\pm$ 0.0 %
Freezersmalltrain Test 73.27 % $\pm$ 5.2 % 75.86 % $\pm$ 0.0 %
Fungi Test 67.74 % $\pm$ 10.2 % 46.24 % $\pm$ 0.0 %
@@
Gesturemidaird1 Test 25.77 % $\pm$ 6.6 % 23.08 % $\pm$ 4.4 %
Gesturemidaird2 Test 30.00 % $\pm$ 7.2 % 21.15 % $\pm$ 0.5 %
Gesturemidaird3 Test 14.81 % $\pm$ 6.9 % 08.46 % $\pm$ 2.2 %
Gesturepebblez1 Test 25.44 % $\pm$ 11.9 % 45.64 % $\pm$ 7.0 %
Gesturepebblez2 Test 27.37 % $\pm$ 13.6 % 26.27 % $\pm$ 6.7 %
Gunpointagespan Test 52.53 % $\pm$ 3.8 % 54.43 % $\pm$ 4.9 %
Gunpointmaleversusfemale Test 55.06 % $\pm$ 5.1 % 84.49 % $\pm$ 0.4 %
Gunpointoldversusyoung Test 52.38 % $\pm$ 0.0 % 100.00 % $\pm$ 0.0 %
Gunpoint Test 54.50 % $\pm$ 7.5 % 57.00 % $\pm$ 10.8 %
Ham Test 58.81 % $\pm$ 6.7 % 51.43 % $\pm$ 0.0 %
Handoutlines Test 64.05 % $\pm$ 0.0 % 64.05 % $\pm$ 0.0 %
Haptics Test 21.19 % $\pm$ 0.8 % 20.78 % $\pm$ 0.0 %
Herring Test 59.38 % $\pm$ 0.0 % 59.38 % $\pm$ 0.0 %
Housetwenty Test 62.18 % $\pm$ 5.9 % 71.01 % $\pm$ 0.6 %
Inlineskate Test 17.68 % $\pm$ 0.7 % 18.00 % $\pm$ 1.0 %
Insectepgregulartrain Test 85.84 % $\pm$ 25.7 % 100.00 % $\pm$ 0.0 %
Insectepgsmalltrain Test 76.00 % $\pm$ 24.5 % 100.00 % $\pm$ 0.0 %
Insectwingbeatsound Test 34.20 % $\pm$ 8.1 % 18.89 % $\pm$ 3.4 %
Italypowerdemand Test 62.05 % $\pm$ 17.6 % 49.85 % $\pm$ 0.0 %
Largekitchenappliances Test 47.27 % $\pm$ 7.8 % 57.73 % $\pm$ 0.2 %
Lightning2 Test 54.10 % $\pm$ 0.0 % 54.10 % $\pm$ 0.0 %
Lightning7 Test 37.67 % $\pm$ 7.3 % 36.99 % $\pm$ 5.8 %
Mallat Test 15.63 % $\pm$ 6.6 % 12.32 % $\pm$ 0.0 %
Meat Test 67.50 % $\pm$ 1.7 % 55.83 % $\pm$ 3.5 %
Medicalimages Test 51.51 % $\pm$ 0.1 % 51.45 % $\pm$ 0.0 %
Melbournepedestrian Test 29.81 % $\pm$ 7.2 % 62.63 % $\pm$ 0.8 %
Middlephalanxoutlineagegroup Test 34.09 % $\pm$ 17.9 % 18.83 % $\pm$ 0.0 %
Middlephalanxoutlinecorrect Test 57.04 % $\pm$ 0.0 % 57.04 % $\pm$ 0.0 %
Middlephalanxtw Test 38.15 % $\pm$ 12.0 % 27.27 % $\pm$ 0.0 %
Mixedshapesregulartrain Test 48.98 % $\pm$ 14.4 % 50.85 % $\pm$ 1.9 %
Mixedshapessmalltrain Test 44.85 % $\pm$ 7.1 % 49.20 % $\pm$ 1.0 %
Motestrain Test 73.06 % $\pm$ 11.4 % 71.45 % $\pm$ 8.3 %
Noninvasivefetalecgthorax1 Test 06.78 % $\pm$ 9.9 % 01.83 % $\pm$ 0.0 %
Noninvasivefetalecgthorax2 Test 08.12 % $\pm$ 12.6 % 02.29 % $\pm$ 0.6 %
Osuleaf Test 29.65 % $\pm$ 8.1 % 18.18 % $\pm$ 0.0 %
Oliveoil Test 40.00 % $\pm$ 0.0 % 40.00 % $\pm$ 0.0 %
Plaid Test 23.32 % $\pm$ 6.1 % 22.81 % $\pm$ 2.2 %
Phalangesoutlinescorrect Test 61.31 % $\pm$ 0.0 % 61.31 % $\pm$ 0.0 %
Phoneme Test 13.13 % $\pm$ 2.0 % 11.29 % $\pm$ 0.0 %
Pickupgesturewiimotez Test 40.00 % $\pm$ 6.7 % 54.00 % $\pm$ 2.8 %
Pigairwaypressure Test 06.61 % $\pm$ 1.7 % 19.71 % $\pm$ 0.0 %
Pigartpressure Test 06.01 % $\pm$ 1.0 % 25.24 % $\pm$ 1.0 %
Pigcvp Test 10.58 % $\pm$ 8.7 % 47.60 % $\pm$ 1.4 %
Plane Test 26.67 % $\pm$ 16.7 % 09.52 % $\pm$ 0.0 %
Powercons Test 88.06 % $\pm$ 3.2 % 85.00 % $\pm$ 0.8 %
Proximalphalanxoutlineagegroup Test 66.22 % $\pm$ 19.9 % 48.78 % $\pm$ 0.0 %
Proximalphalanxoutlinecorrect Test 68.38 % $\pm$ 0.0 % 68.38 % $\pm$ 0.0 %
Proximalphalanxtw Test 50.49 % $\pm$ 17.8 % 35.12 % $\pm$ 0.0 %
Refrigerationdevices Test 41.20 % $\pm$ 4.5 % 52.00 % $\pm$ 3.4 %
Rock Test 44.50 % $\pm$ 3.4 % 39.00 % $\pm$ 7.1 %
@@
Screentype Test 37.33 % $\pm$ 5.4 % 41.33 % $\pm$ 0.8 %
Semghandgenderch2 Test 68.29 % $\pm$ 5.8 % 55.25 % $\pm$ 5.1 %
Semghandmovementch2 Test 27.44 % $\pm$ 7.0 % 32.00 % $\pm$ 4.4 %
Semghandsubjectch2 Test 40.67 % $\pm$ 8.7 % 33.67 % $\pm$ 2.7 %
Shakegesturewiimotez Test 46.00 % $\pm$ 6.3 % 71.00 % $\pm$ 9.9 %
Shapeletsim Test 60.83 % $\pm$ 8.7 % 83.33 % $\pm$ 6.3 %
Shapesall Test 29.46 % $\pm$ 1.0 % 43.00 % $\pm$ 0.5 %
Smallkitchenappliances Test 48.60 % $\pm$ 9.5 % 55.33 % $\pm$ 0.2 %
Smoothsubspace Test 61.17 % $\pm$ 5.3 % 77.33 % $\pm$ 0.9 %
Sonyaiborobotsurface1 Test 47.67 % $\pm$ 8.1 % 42.93 % $\pm$ 0.0 %
Sonyaiborobotsurface2 Test 67.24 % $\pm$ 7.7 % 61.70 % $\pm$ 0.0 %
Starlightcurves Test 76.83 % $\pm$ 12.8 % 83.96 % $\pm$ 1.1 %
Strawberry Test 64.32 % $\pm$ 0.0 % 64.32 % $\pm$ 0.0 %
Swedishleaf Test 25.72 % $\pm$ 6.0 % 07.84 % $\pm$ 0.9 %
Symbols Test 43.52 % $\pm$ 20.0 % 17.39 % $\pm$ 0.0 %
Syntheticcontrol Test 62.67 % $\pm$ 14.2 % 82.17 % $\pm$ 6.8 %
Toesegmentation1 Test 62.17 % $\pm$ 5.9 % 66.89 % $\pm$ 1.6 %
Toesegmentation2 Test 75.77 % $\pm$ 11.8 % 73.08 % $\pm$ 10.9 %
Trace Test 52.25 % $\pm$ 11.1 % 19.00 % $\pm$ 0.0 %
Twoleadecg Test 51.51 % $\pm$ 3.1 % 49.96 % $\pm$ 0.0 %
Twopatterns Test 34.88 % $\pm$ 4.1 % 25.87 % $\pm$ 0.0 %
Umd Test 52.78 % $\pm$ 5.9 % 55.90 % $\pm$ 8.3 %
Uwavegesturelibraryall Test 26.95 % $\pm$ 18.5 % 12.33 % $\pm$ 0.1 %
Uwavegesturelibraryx Test 57.49 % $\pm$ 4.6 % 17.23 % $\pm$ 7.1 %
Uwavegesturelibraryy Test 48.75 % $\pm$ 6.1 % 19.11 % $\pm$ 1.8 %
Uwavegesturelibraryz Test 51.70 % $\pm$ 5.4 % 25.49 % $\pm$ 8.3 %
Wafer Test 90.97 % $\pm$ 3.5 % 89.21 % $\pm$ 0.0 %
Wine Test 50.00 % $\pm$ 0.0 % 50.00 % $\pm$ 0.0 %
Wordsynonyms Test 27.35 % $\pm$ 4.5 % 21.94 % $\pm$ 0.0 %
Wormstwoclass Test 58.44 % $\pm$ 2.6 % 57.14 % $\pm$ 0.0 %
Worms Test 43.83 % $\pm$ 1.9 % 42.86 % $\pm$ 0.0 %
Yoga Test 53.34 % $\pm$ 0.3 % 53.57 % $\pm$ 0.0 %